# On a novel UCP result and its application to inverse conductive scattering

### v3

@@ 1.3. Connections to existing results and main contributions
In what follows, we shall prove this theorem by using mathematical induction. Namely, we shall prove that $a_{\ell}=b_{\ell}=0$ for $\ell\in\mathbb{N}$. Since $\gamma_{2}\in H^{2}(\mathcal{S}_{r_{0}})$, let $\widetilde{\gamma_{2}}$ be the Sobolev extension of $\gamma_{2}$ in $\mathbb{R}^{2}$. According to Lemma 2.1, there exists the CGO solution $u_{0}$ given by (2.1), which satisfies
@@
$A=\left(\begin{matrix}e^{-\mathrm{i}\vartheta_{m}}+e^{-\mathrm{i}\vartheta_{M}}&e^{-3\mathrm{i}\vartheta_{m}}+e^{-3\mathrm{i}\vartheta_{M}}\\ e^{3\mathrm{i}\vartheta_{m}}+e^{-3\mathrm{i}\vartheta_{M}}&e^{\mathrm{i}\vartheta_{m}}+e^{\mathrm{i}\vartheta_{M}}\end{matrix}\right).$
@@
In the following we will prove Theorem 1.2.
The proof of Theorem 1.2.
We prove this theorem using a contradiction argument. Assume that $\Omega$ has an irrational convex corner and is invisible. Since the Laplacian $\Delta$ is invariant under rigid motion, this corner of $\Omega$ can be described by $\mathcal{S}_{r_{0}}$ as given by (1.2). Since $u^{\infty}\equiv 0$, by Rellich’s theorem, one has the PDE system (1.3), where $v=u^{i}$, $w=u^{-}\chi_{\Omega}+u^{i}\chi_{D\backslash\overline{\Omega}}$, $\gamma_{1}=k^{2}$ and $\gamma_{2}=k^{2}q\chi_{\Omega}+k^{2}\chi_{D\backslash\overline{\Omega}}$. Here $D:=B_{R}$ is chosen such that $\Omega\Subset D$. By applying Theorem 1.1, we conclude that $u^{i}\equiv 0$ in $D$, leading to a contradiction by unique continuation principle. ∎
To address the inverse problem (1.10), it is essential to define the admissible class of scatterers. Definition 2.1 introduces the concept of an irrational polygon, which will be used in Definition 2.2.
Definition 2.1.
Suppose $\Omega\subset\mathbb{R}^{2}$ is a polygon. Let $\omega=\lambda\cdot\pi$, where $\lambda\in(0,2)$, be an interior angle of $\Omega$. An angle $\omega$ is considered irrational if $\lambda$ is an irrational number; otherwise, it is considered rational. Furthermore, if $\lambda\in(0,1)$ is an irrational number, then the angle $\omega$ is considered convex irrational. The polygon $\Omega$ is termed an irrational polygon if each of its interior angles is an irrational angle; otherwise, it is considered rational.
@@
$(\Omega;q,\eta)=\bigcup_{\ell=1}^{N}(\Omega_{\ell};q_{\ell},\eta_{\ell}),\ \Omega_{\ell}\cap\Omega_{m}=\emptyset\ \mbox{for}\ \ell\not=m,$
@@
The following theorem provides a local uniqueness result for determining the shape of a scatterer through a single far-field measurement, offering a more rigorous version of Theorem 1.3. It asserts that the difference set between two admissible complex polygonal scatterers cannot contain a convex irrational corner if the far-field measurements associated with these two scatterers, for a fixed incident wave, are identical. Furthermore, we establish the global unique identification of the shape of admissible complex convex irrational polygonal scatterers using only a single far-field measurement.
@@
$(\Omega_{\ell};q_{\ell},\eta_{\ell})=\bigcup_{i=1}^{N_{\ell}}(\Omega_{i,\ell};q_{i,\ell},\eta_{i,\ell}),\quad N_{\ell}\in\mathbb{N}{+},\quad\ell=1,2.$
@@
for all $\mathbf{\hat{x}}\in\mathbb{S}^{1}$ and a fixed incident wave $u^{i}$, then $\Omega_{1}\Delta\Omega_{2}=(\Omega_{1}\backslash\Omega_{2})\cup(\Omega_{2}\backslash\Omega_{1})$ cannot contain a convex irrational corner.
@@
We prove the first conclusion of this theorem by contradiction. When (2.18) holds, assume that $\Omega_{1}\Delta\Omega_{2}$ contains an irrational corner. Underlying the fact that $\Delta$ is variant under the rigid motion and by the definition of an admissible complex irrational polygonal scatterer, we also assume that there exists an irrational corner $\mathcal{S}_{r_{0}}$ such that $\mathcal{S}_{r_{0}}\Subset\Omega_{1}\setminus\overline{\Omega_{2}}$, where $r_{0}\in\mathbb{R}_{+}$ is sufficiently small. Let $B_{r_{0}}$ be a disk with $r_{0}\in\mathbb{R}_{+}$ be chosen such that $B_{r_{0}}\Subset\Omega_{1}\setminus\overline{\Omega_{2}}$ and $\mathcal{S}_{r_{0}}\subset B_{r_{0}}$.
@@ 3. Unique recovery results for the conductive polygonal-nest and polygonal-cell medium scatterers
In order to introduce the admissible class for studying the inverse problems of the unique identification of the shape, we need to study the Hölder-regularity of the solution of the coupled PDEs system (3.7) near the corner, where $u_{1}$ and $u_{2}$ are the total wave field corresponding to the two different scatterers $\Omega_{1}$ and $\Omega_{2}$, respectively. Indeed, the PDEs system (3.7) is derived from the proceed of proving the unique determination of the shape of the scatterer by contradiction (see (3.12) in the proof of Theorem 3.1).
@@
where $k$ is a positive constant and $q$ is in the form of (3.1). If $u_{2}$ is real analytic in $\mathcal{S}_{2r_{0}}$, then $u_{1}\in C^{1,\alpha}(\overline{\mathcal{S}_{r_{0}}})$, where $\alpha\in(0,1)$.
Before proving Lemma 3.1, we first introduce a triangle $V_{2r_{0}}$ described below, which will be used in a subsequent argument. Let us pick up the points $\mathbf{y}_{+}$ and $\mathbf{y}_{-}$ on $\Gamma_{2r_{0}}^{+}$ and $\Gamma_{2r_{0}}^{-}$, respectively, where $|\mathbf{y}_{\pm}|\in(r_{0},2r_{0})$. Denote $\Lambda_{2r_{0}}^{{}^{\prime}}=\mathbf{y}_{+}-\mathbf{y}_{-}$. Then the triangle domain $V_{2r_{0}}$ is formed by three vectors $\mathbf{y}_{+}$, $\mathbf{y}_{-}$ and $\Lambda_{2r_{0}}^{{}^{\prime}}$. It is clear that $\overline{\mathcal{S}_{r_{0}}}\subset\overline{V_{2r_{0}}}\subset\overline{\mathcal{S}_{2r_{0}}}$. We can refer to Fig. 2 for a geometric illustration of $V_{2r_{0}}$.
Figure 2. A schematic illustration of $\mathcal{S}_{2r_{0}}$, $V_{2r_{0}}$ and $\mathcal{S}_{r_{0}}$.
The proof of Lemma 3.1.
Let $w=u_{1}-u_{2},$ then from (3.7) it can be directly calculated that
$\begin{cases}\Delta w=f,\hskip 15.6491pt\mbox{in}\ V_{2r_{0}},\\ w=\varphi,\hskip 22.76228pt\mbox{on}\ \Lambda^{{}^{\prime}}_{2r_{0}},\\ w=0,\hskip 24.18501pt\mbox{on}\ \Gamma_{2r_{0}}^{\pm},\end{cases}$
[…]
$w({\bf x})=w_{\rm reg}({\bf x})+\sum_{0<\lambda_{\ell}<3-\frac{2}{p}}\zeta^{\ell}r^{\lambda_{\ell}}\sin(\lambda_{\ell}\vartheta),\ w_{\rm reg}\in W^{3,p}(V_{2r_{0}}),\ {\bf x}=r(\cos\vartheta,\sin\vartheta),$
where $\zeta$ is a constant depending only on $f$ and $\varphi$, and $\lambda_{\ell}=\frac{\ell\pi}{\vartheta_{M}-\vartheta_{m}},\ \ell\in\mathbb{N}$. If $\vartheta_{M}-\vartheta_{m}\in(0,\pi/2]$, then we know that $w({\bf x})=w_{\rm reg}({\bf x})$. If $\vartheta_{M}-\vartheta_{m}\in(\pi/2,\pi)$, it can be verified that $\frac{2}{3-\frac{\pi}{\vartheta_{M}-\vartheta_{m}}}\in(1,2)$. Hence we can choose $p\in(1,\frac{2}{3-\frac{\pi}{\vartheta_{M}-\vartheta_{m}}})$. This indicates that $\lambda_{\ell}>3-\frac{2}{p}$ for any $\ell\in\mathbb{N}$. Therefore, $w=w_{\rm reg}\in W^{3,p}(V_{2r_{0}})$. By virtue of the Sobolev embedding theorem (cf. [28, Theorem 7.26]), we have that $w=w_{\rm reg}\in C^{1,\alpha}(\overline{V_{2r_{0}}})$ for $\alpha\in(0,1)$.
@@
Suppose $(\Omega;q,\eta)$ represents a conductive polygonal-nest or polygonal-cell medium scatterer as defined in Definitions 3.1 or 3.2. Let $u$ denote the total wave field of (1.7) associated with $(\Omega;q,\eta)$.
@@
        $\lim_{\rho\to 0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|u(\mathbf{x}_{p})|\not=0,$ (3.8)
@@
        $\lim_{\rho\to 0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|u(\mathbf{x}_{p})|\mathrm{d}\mathbf{x}=0,\ \mbox{then}$
        $\lim_{\rho\to 0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|\nabla u(\mathbf{x}_{p})|\mathrm{d}\mathbf{x}\neq 0,$ (3.9)
@@
where $C$ is a positive number and $B_{R}$ is a ball centered at the origin with radius $R$. Hence, if (3.10) is satisfied and the conductive boundary parameter $\eta$ is sufficiently small, the plane incident wave $u^{i}$ dominates the total wave $u=u^{i}+u^{s}$. It is evident that $u^{i}$ fulfills (3.8) (or (3.9)) , which implies that the total wave $u$ generically satisfies (3.8) or (3.9). On the other hand, from a physical standpoint, in conductive medium scattering, if the total wave does not satisfies (3.8) at a corner point, it implies that the scattered wave and the incident wave cancel each other at those points. Similarly, if the admissible condition (3.9) is violated, it indicates that either the scattered wave and the incident wave cancel each other, or the gradient of the scattered wave and the incident wave cancel each other at those points. Such cancellation phenomena, as described above, are rare in typical scattering scenarios. However, we will not delve deeply into this matter in this paper.
@@
    $\lim_{\rho\to 0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|\partial_{1}u(\mathbf{x}_{p})|\mathrm{d}\mathbf{x}\neq 0$,
    $\lim_{\rho\to 0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|u(\mathbf{x}_{p})|\mathrm{d}\mathbf{x}=\lim_{\rho\to 0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|\partial_{2}u(\mathbf{x}_{p})|\mathrm{d}\mathbf{x}=0$,
@@
    $\lim_{\rho\to 0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|\partial_{2}u(\mathbf{x}_{p})|\mathrm{d}\mathbf{x}\neq 0$,
    $\lim_{\rho\to 0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|u(\mathbf{x}_{p})|\mathrm{d}\mathbf{x}=\lim_{\rho\to 0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|\partial_{1}u(\mathbf{x}_{p})|\mathrm{d}\mathbf{x}=0$,
@@
    $\lim_{\rho\to 0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|\partial_{1}u(\mathbf{x}_{p})|\mathrm{d}\mathbf{x}\neq 0$,
    $\lim_{\rho\to 0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|\partial_{2}u(\mathbf{x}_{p})|\mathrm{d}\mathbf{x}\neq 0$, $\lim_{\rho\to 0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|u(\mathbf{x}_{p})|\mathrm{d}\mathbf{x}=0$.
@@
Let $(\Omega_{\ell};q_{\ell},\eta_{\ell}),\ell=1,2$ be two admissible conductive polygonal-nest or polygonal-cell medium scatterer satisfying (1.7), $u^{\infty}_{\ell}(\mathbf{\hat{x}};u^{i})$ be the far-field pattern associated with the scatterer $(\Omega_{\ell};q_{\ell},\eta_{\ell})$ and the incident field $u^{i}$. If
@@
is fulfilled for all $\mathbf{\hat{x}}\in\mathbb{S}^{1}$ and a fixed incident wave $u^{i}$, then $\Omega_{1}\Delta\Omega_{2}$ cannot have a convex corner. Moreover, if $\Omega_{1}$ and $\Omega_{2}$ are two admissible conductive polygonal-nest medium scatterers, one holds that
@@
We only present the proof for the polygonal-nest case, since the polygonal-cell case can be proven in a similar manner. We proceed to prove this theorem by contradiction. Let us assume that there exists a corner in $\Omega_{1}\Delta\Omega_{2}$. According to (3.2), we can express the admissible conductive polygonal-nest medium scatterers as $(\Omega_{\ell};q_{\ell},\eta_{\ell})$ ($\ell=1,2$) as follows
@@
According to Lemma 3.1, we know that $u_{1}^{-}\in C^{1,\alpha}(\overline{\mathcal{S}_{r_{0}}}),\alpha\in(0,1)$, which implies the following expansion

### v4

@@ 1.3. Connections to existing results and main contributions
In what follows, we shall prove this theorem by using mathematical induction. Namely, we shall prove that $a_{\ell}=b_{\ell}=0$ for $\ell\in\mathbb{N}$. Since $\gamma_{2}\in H^{2}(\mathcal{S}_{r_{0}})$, let $\widetilde{\gamma_{2}}$ be the Sobolev extension of $\gamma_{2}$ in $H^{2}(\mathbb{R}^{2})$. According to Lemma 2.1, there exists the CGO solution $u_{0}$ given by (2.1), which satisfies
@@
$A=\left(\begin{matrix}e^{-\mathrm{i}\vartheta_{m}}+e^{-\mathrm{i}\vartheta_{M}}&e^{-3\mathrm{i}\vartheta_{m}}+e^{-3\mathrm{i}\vartheta_{M}}\\ e^{3\mathrm{i}\vartheta_{m}}+e^{3\mathrm{i}\vartheta_{M}}&e^{\mathrm{i}\vartheta_{m}}+e^{\mathrm{i}\vartheta_{M}}\end{matrix}\right).$
@@
In the following we will prove Corollary 1.2.
The proof of Corollary 1.2.
We prove this theorem using a contradiction argument. Assume that $\Omega$ has an irrational convex corner and is invisible. Since the Laplacian $\Delta$ is invariant under rigid motion, this corner of $\Omega$ can be described by $\mathcal{S}_{r_{0}}$ as given by (1.3). Since $u^{\infty}\equiv 0$, by Rellich’s theorem, one has the PDE system (1.4), where $v=u^{i}$, $w=u^{-}\chi_{\Omega}+u^{i}\chi_{D\backslash\overline{\Omega}}$, $\gamma_{1}=k^{2}$ and $\gamma_{2}=k^{2}q\chi_{\Omega}+k^{2}\chi_{D\backslash\overline{\Omega}}$. Here $D:=B_{R}$ is chosen such that $\Omega\Subset D$. By applying Theorem 1.1, we conclude that $u^{i}\equiv 0$ in $D$, which leads to a contradiction by the unique continuation principle. ∎
To address the inverse problem (1.9), it is essential to define the admissible class of scatterers. Recall that Definition 1.1 introduces the concept of an irrational polygon, which will be utilized in Definition 2.1.
@@
$(\Omega;q,\eta)=\bigcup_{\ell=1}^{N}(\Omega_{\ell};q_{\ell},\eta_{\ell}),\ \overline{\Omega_{\ell}}\cap\overline{\Omega_{m}}=\emptyset\ \mbox{for}\ \ell\not=m,$
@@
The following theorem provides a local uniqueness result for determining the shape of a scatterer through a single far-field measurement, offering a more rigorous version of Corollary 1.3. It asserts that the difference set between two admissible complex polygonal scatterers cannot contain a convex irrational corner if the far-field measurements associated with these two scatterers, for a fixed incident wave, are identical. Furthermore, we establish the global unique identification of the shape of admissible complex convex irrational polygonal scatterers using only a single far-field measurement.
@@
$(\Omega_{\ell};q_{\ell},\eta_{\ell})=\bigcup_{i=1}^{N_{\ell}}(\Omega_{i,\ell};q_{i,\ell},\eta_{i,\ell}),\quad N_{\ell}\in\mathbb{N}_{+},\quad\ell=1,2.$
@@
for all $\mathbf{\hat{x}}\in\mathbb{S}^{1}$ and a fixed incident wave $u^{i}$, then the difference $\Omega_{1}\aleph\Omega_{2}$ defined in Corollary 1.3 cannot contain a convex irrational corner.
@@
We prove the first conclusion of this theorem by contradiction. When (2.18) holds, assume that $\Omega_{1}\aleph\Omega_{2}$ contains an irrational corner. Underlying the fact that the Laplacian $\Delta$ is invariant under the rigid motion and by the definition of an admissible complex irrational polygonal scatterer, we also assume that there exists an irrational corner $\mathcal{S}_{r_{0}}$ such that $\mathcal{S}_{r_{0}}\Subset\Omega_{1}\setminus\overline{\Omega_{2}}$, where $r_{0}\in\mathbb{R}_{+}$ is sufficiently small. Let $B_{r_{0}}$ be a disk with radius $r_{0}\in\mathbb{R}_{+}$.
@@ 3. Unique recovery results for the conductive polygonal-nest and polygonal-cell medium scatterers
In order to introduce the admissible class for studying the inverse problems of the unique identification of the shape, we need to study the Hölder-regularity of the solution of the coupled PDE system (3.7) near the corner, where $u_{1}$ and $u_{2}$ are the total wave field corresponding to the two different scatterers $\Omega_{1}$ and $\Omega_{2}$, respectively. Indeed, the PDE system (3.7) is derived from the proceed of proving the unique determination of the shape of the scatterer by contradiction (see (3.13) in the proof of Theorem 3.1). Recall that $\mathcal{S}_{r_{0}}$ is defined by (1.3) with two line segment boundaries $\Gamma^{\pm}_{r_{0}}$. We can similarly define $\mathcal{S}_{tr_{0}}$ for $t\in\mathbb{R}_{+}$.
@@
$\begin{cases}(\Delta+k^{2}q)u_{1}=0\hskip 28.45274pt\mbox{in}\ \mathcal{S}_{2r_{0}},\\ (\Delta+k^{2})u_{2}=0\hskip 34.14322pt\mbox{in}\ B_{2r_{0}},\\ u_{1}=u_{2}\hskip 68.28644pt\mbox{on}\ \Gamma_{2r_{0}}^{\pm},\end{cases}$ (3.7)
where $k$ is a positive constant and $q$ is in the form of (3.1). There exists a positive number $t_{0}^{\prime}\in\mathbb{R}_{+}$ such that $u_{1}\in C^{1,\alpha}(\overline{\mathcal{S}_{t_{0}^{\prime}r_{0}}})$, where $\alpha\in(0,1)$ depends on the opening angle of $\mathcal{S}_{2r_{0}}$.
Proof.
For any given $t\in\mathbb{R}_{+}$, let $\mathbf{y}_{t,+}\in\Gamma^{+}$ and $\mathbf{y}_{t,-}\in\Gamma^{-}$ be points satisfying $|\mathbf{y}_{t,+}|=|\mathbf{y}_{t,-}|=t$, where $\Gamma^{+}$ and $\Gamma^{-}$ are defined by (1.2a) and (1.2b), respectively. For any $t\in\mathbb{R}_{+}$, we introduce the triangle $V_{tr_{0}}$ formed by the three line segments connecting $\mathbf{0}$, $\mathbf{y}_{t,+}$, and $\mathbf{y}_{t,-}$.
Let $\varphi\in C^{\infty}({\mathcal{S}_{2r_{0}}})$ be a smooth cutoff function satisfying that $\varphi\equiv 1$ in ${\mathcal{S}_{t_{0}r_{0}}}$ and $\varphi\equiv 0$ in $\mathcal{S}_{2r_{0}}\setminus\overline{\mathcal{S}_{t_{0}r_{0}}}$, where $t_{0}\in(0,2)$ such that $\mathcal{S}_{t_{0}r_{0}}\Subset V_{2r_{0}}$. Let $w=u_{1}-u_{2}$ and define $\tilde{\varphi}=w\varphi$. Then there holds that
$\Delta\tilde{\varphi}=f\ \mbox{in}\ V_{2r_{0}},\quad\tilde{\varphi}=0\ \mbox{on}\ \partial{V_{2{r_{0}}}},$
where $f=\Delta w\varphi+w\Delta\varphi+2\nabla w\cdot\nabla\varphi=(k^{2}u_{2}-k^{2}qu_{1})\varphi+w\Delta\varphi+2\nabla w\cdot\nabla\varphi.$ Since $u_{1}\in H^{1}(\mathcal{S}_{2r_{0}})$ and $u_{2}$ is analytic in $\mathcal{S}_{2r_{0}}$, it yields that $w\in H^{1}(\mathcal{S}_{2r_{0}})$, $\nabla w\in L^{2}(\mathcal{S}_{2r_{0}})$ and $f\in L^{2}(\mathcal{S}_{2r_{0}})$. Moreover, it is noted that the opening angle of the sector $\vartheta_{M}-\vartheta_{m}\in(0,\pi)$ and $\tilde{\varphi}\in H^{1}_{0}(V_{2r_{0}})$. Applying [22, Theorem 10.28] to the case of the polygonal network consists of only one subset $V_{2r_{0}}$, then we have $\tilde{\varphi}\in H^{2}(V_{2r_{0}})$. Consequently, we have $w\in H^{2}(\mathcal{S}_{t_{0}r_{0}})$ since $\varphi\equiv 1$ in $\mathcal{S}_{t_{0}r_{0}}$. Therefore, we conclude that $u_{1}\in H^{2}(\mathcal{S}_{t_{0}r_{0}})$.
Now we focus on the triangular domain $V_{t_{0}r_{0}}$. Let $\psi\in C^{\infty}(\mathcal{S}_{t_{0}r_{0}})$ be a smooth cutoff function satisfying:
$\psi\equiv 1\text{ in }\mathcal{S}_{t_{0}^{\prime}r_{0}}\quad\text{and}\quad\psi\equiv 0\text{ in }\mathcal{S}_{t_{0}r_{0}}\setminus\mathcal{S}_{t_{0}^{\prime}r_{0}},$
where $t_{0}^{\prime}$ is fixed such that $0<t_{0}^{\prime}<t_{0}$ and $\mathcal{S}_{t_{0}^{\prime}r_{0}}\Subset V_{t_{0}r_{0}}$. Similarly, denote $\tilde{\psi}=w\psi$, we then have
$\Delta\tilde{\psi}=g\ \mbox{in}\ V_{t_{0}r_{0}},\quad\tilde{\psi}=0\ \mbox{on}\ \partial V_{t_{0}r_{0}},$ (3.8)
@@
$\tilde{\psi}({\bf x})=\tilde{\psi}_{\rm reg}({\bf x})+\sum_{0<\lambda_{\ell}<3-\frac{2}{p}}\zeta_{\ell}r^{\lambda_{\ell}}\sin(\lambda_{\ell}\vartheta),\ \tilde{\psi}_{\rm reg}\in W^{3,p}(V_{2r_{0}}),\ {\bf x}=r(\cos\vartheta,\sin\vartheta),$
where $\zeta$ is a constant depending only on $f$ and $\varphi$, and $\lambda_{\ell}=\frac{\ell\pi}{\vartheta_{M}-\vartheta_{m}},\ \ell\in\mathbb{N}_{+}$. If $\vartheta_{M}-\vartheta_{m}\in(0,\pi/2]$, then we know that $\tilde{\psi}({\bf x})=\tilde{\psi}_{\rm reg}({\bf x})$ since $\lambda_{\ell}\geq 2\ell>3-\frac{2}{p}$. If $\vartheta_{M}-\vartheta_{m}\in(\pi/2,\pi)$, we then choose $p\in(1,\frac{2}{3-\frac{\pi}{\vartheta_{M}-\vartheta_{m}}})$ since it can be verified that $\frac{2}{3-\frac{\pi}{\vartheta_{M}-\vartheta_{m}}}\in(1,2)$, which also indicates that $\lambda_{\ell}>3-\frac{2}{p}$ for any $\ell\in\mathbb{N}_{+}$. Therefore, $\tilde{\psi}=\tilde{\psi}_{\rm reg}\in W^{3,p}(V_{t_{0}r_{0}})$ for $p\in(1,2)$. Recall that $\psi\equiv 1$ in $\mathcal{S}_{t_{0}^{\prime}r_{0}}$, then we deduce that $w\in W^{3,p}(\mathcal{S}_{t_{0}^{\prime}r_{0}})$ which implies that $u_{1}\in W^{3,p}(\mathcal{S}_{t_{0}^{\prime}r_{0}})$ since $u_{2}$ is analytic in $\mathcal{S}_{t_{0}^{\prime}r_{0}}$. It follows from the Sobolev embedding theorem (cf. [38, Theorem 7.26]), we have that $u_{1}\in C^{1,\alpha}(\overline{\mathcal{S}_{t_{0}^{\prime}r_{0}}})$ with $\alpha\in(0,1)$, where $\alpha$ depends on the opening angle of the sector $\mathcal{S}_{2r_{0}}$.
@@
Suppose $(\Omega;q,\eta)$ represents a conductive polygonal-nest or polygonal-cell medium scatterer as defined in Definitions 3.3 or 3.4. Let $u$ denote the total wave field of (1.6) associated with $(\Omega;q,\eta)$.
@@
        $\lim_{\rho\to+0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|u(\mathbf{x})|\mathrm{d}\mathbf{x}\not=0,$ (3.9)
@@
        $\lim_{\rho\to+0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|u(\mathbf{x})|\mathrm{d}\mathbf{x}=0,\ \mbox{then}$
        $\lim_{\rho\to+0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|\nabla u(\mathbf{x})|\mathrm{d}\mathbf{x}\neq 0,$ (3.10)
@@
where $C$ is a positive number and $B_{R}$ is a ball centered at the origin with radius $R$. Hence, if (3.11) is satisfied and the conductive boundary parameter $\eta$ is sufficiently small, the incident wave $u^{i}$ dominates the total wave $u=u^{i}+u^{s}$. It is evident that the incident wave $u^{i}$ fulfills (3.9) (or (3.10)) , which implies that the total wave $u$ generically satisfies (3.9) or (3.10). On the other hand, from a physical standpoint, in conductive medium scattering, if the total wave does not satisfies (3.9) at a corner point, it implies that the scattered wave and the incident wave cancel each other at those points. Similarly, if the admissible condition (3.10) is violated, it indicates that either the scattered wave and the incident wave cancel each other, or the gradient of the scattered wave and the incident wave cancel each other at those points. Such cancellation phenomena, as described above, are rare in typical scattering scenarios. However, we will not delve deeply into this matter in this paper.
@@
    $\lim_{\rho\to+0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|\partial_{1}u(\mathbf{x})|\mathrm{d}\mathbf{x}\neq 0$,
    $\lim_{\rho\to+0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|u(\mathbf{x})|\mathrm{d}\mathbf{x}=\lim_{\rho\to+0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|\partial_{2}u(\mathbf{x})|\mathrm{d}\mathbf{x}=0$,
@@
    $\lim_{\rho\to+0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|\partial_{2}u(\mathbf{x})|\mathrm{d}\mathbf{x}\neq 0$,
    $\lim_{\rho\to+0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|u(\mathbf{x})|\mathrm{d}\mathbf{x}=\lim_{\rho\to+0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|\partial_{1}u(\mathbf{x})|\mathrm{d}\mathbf{x}=0$,
@@
    $\lim_{\rho\to+0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|\partial_{1}u(\mathbf{x})|\mathrm{d}\mathbf{x}\neq 0$,
    $\lim_{\rho\to+0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|\partial_{2}u(\mathbf{x})|\mathrm{d}\mathbf{x}\neq 0$, $\lim_{\rho\to+0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|u(\mathbf{x})|\mathrm{d}\mathbf{x}=0$.
@@
Let $(\Omega_{\ell};q_{\ell},\eta_{\ell}),\ell=1,2$ be both admissible conductive polygonal-nest or both polygonal-cell medium scatterers satisfying (1.6), $u^{\infty}_{\ell}(\mathbf{\hat{x}};u^{i})$ be the far-field pattern associated with the scatterers $(\Omega_{\ell};q_{\ell},\eta_{\ell})$ and the incident field $u^{i}$. If
@@
is fulfilled for all $\mathbf{\hat{x}}\in\mathbb{S}^{1}$ and a fixed incident wave $u^{i}$, then the difference $\Omega_{1}\aleph\Omega_{2}$ defined in Corollary 1.3 cannot have a convex corner. Moreover, if $\Omega_{1}$ and $\Omega_{2}$ are both admissible conductive polygonal-nest medium scatterers, one holds that
@@
We only present the proof for the polygonal-nest case, since the polygonal-cell case can be proven in a similar manner. We proceed to prove this theorem by contradiction. Let us assume that there exists a corner in $\Omega_{1}\aleph\Omega_{2}$. According to (3.2), we can express the admissible conductive polygonal-nest medium scatterers as $(\Omega_{\ell};q_{\ell},\eta_{\ell})$ ($\ell=1,2$) as follows
@@
According to Lemma 3.1, there exists a positive number $t_{0}^{\prime}\in\mathbb{R}_{+}$ such that $u_{1}^{-}\in C^{1,\alpha}(\overline{\mathcal{S}_{t_{0}^{\prime}r_{0}}}),\alpha\in(0,1)$, which implies the following expansion
@@
In the subsequent analysis, we need to establish an integral identity in $\mathcal{S}_{t_{0}^{\prime}r_{0}}$ that is analogous to (3.14). However, our analysis indicates that the parameter $t_{0}^{\prime}$ defining $\mathcal{S}_{t_{0}^{\prime}r_{0}}$ does not play a crucial role in proving this theorem. Therefore, we will instead use (3.14) for our further analysis.